The most recent financial statements for Fleury Inc., follow. Sales for 2012 are projected to grow by 20 percent. Interest expense will remain constant, the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets and accounts payable increase spontaneously with sales FLEURY, INC 2011 Income Statement Sales Costs Other expenses $ 758,000 593,000 14,000 Earnings before interest and taxes Interest paid $151,000 10,000 Taxable income Taxes (40%) $141,000 56,400 Net income $ 84,600 Dividends Addition to retained earnings $ 33,840 50,760 FLEURY, INC Balance Sheet as of December 31, 2011 Assets Liabilities and Owners Equity Current assets Current liabilities Cash Accounts receivable S 21,740 Accounts payable $ 55,900 34,060 Notes payable 15,100 $ 71,000 $ 102,000 Common stock and paid-in surplus $102,000 71,020 Total Inventory Total Fixed assets S 126,820 Long-term debt Owners equity Net plant and equipment S275,000 126,820 $ 228,820 401,820 Total liabilities and owners equity 401,820 Retained earnings Total Total assets What is the EFN if the firm wishes to keep its debt-equity ratio constant? (Do not round intermediate calculations.) EFN
Answer #1

Answer:

Projected income statement for 2012 is given below:

FLEURY INC Income Statement 2011 Projected for 2012 $909,600 $711,600 $16,800 $181,200 $10,000 $171,200 $68,480 $102,720 Calc

Δ Sales in 2012= $758,000 * 20% = $151,600

Existing debt = Long term debt + Current liability = 102,000 + 71,000 =173,000

As the firm wants to keep debt-equity ratio constant, total debt can increase by = (Existing debt / Existing equity) * Projected Addition to retained earnings

= (173000 / 228820) * 61632 = $46,597

Increase in accounts payable = (Existing Accounts payable /Sales) * Δ Sales = (55900/ 758,000) * $151,600 = $11,180

As $46,597 > $11,180

Hence:

If the firm wants to keep debt-equity ratio constant, EFN =

= Total Assets * Δ Sales / Sales - Projected addition to retained earnings - Projected addition required to 'total debt' to keep debt-equity ratio constant

= 401820 * 20% - 61632 - 46597

= - $27,865

$ -27,865